Karen is 6 years older than her sister Michelle, and Michelle is 2 years younger than their brother David. If the sum of their ages is 62 , how old is Michelle?
Answer by ankor@dixie-net.com(22740)

Karen is 6 years older than her sister Michelle,
k = m + 6
:
and Michelle is 2 years younger than their brother David.
m = d - 2
or
d = m + 2
:
If the sum of their ages is 62 , how old is Michelle?
k + m + d = 62
replace k and d
(m+6) + m + (m+2) = 62
3m + 8 = 62
3m = 62 - 8
3m = 54
m = 54/3
m = 18 yrs is M's age
:
:
You can confirm this yourself, find the ages of k and d, check the total